OpenStudy (anonymous):

Harriet needs to ship a small vase. the box she will use has a volume of 216 inches. If the side lengths are all the same what is the length of each side of the box?

OpenStudy (anonymous):

Volume of a box is length * width * height So if the side length is "s", then the volume is s * s * s, right? that means s * s* s = 216

OpenStudy (anonymous):

The answer is 72. Like JakeV8 said, s*s*s=216 so just divide 216 by 3.

OpenStudy (anonymous):

that's not right, actually. s * s * s is "s cubed", not "3 times s" like if s = 2, it would be 2 * 2 * 2 = 8, not 2*3 = 6

OpenStudy (anonymous):

In that case the answer would be six right?

OpenStudy (anonymous):

so the question is really asking "what number cubed is equal 216?" Yes, it's s= 6. If you have a calculator with a cube root button, you can do it that way. You could also try different side lengths and just see which works. 2^3 = 8 3^3 = 27 4^3 = 64 5^3 = 125 6^3 = 216

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Thanks for both of your help!

OpenStudy (anonymous):

Glad to help :) Does it make sense now?

OpenStudy (anonymous):

yes thank you